Find the whole quantity if 
8% of it is 40 liters
Find the whole quantity if 
8% of it is 40 liters

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the whole quantity be x, Then,
(8/100) × (x) = 40
X = 40 × (100/8)
X = 40 × (100/8)
X = 40 × 12.5
X = 500 liters

Thanks